OpenStudy (anonymous):

How to do (3,-2); m=1/2

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (anonymous):

x1 and y1 are the x and y values of your point

OpenStudy (anonymous):

so its 2-(-2)/1-3

OpenStudy (anonymous):

umm, well i assume you're trying to find the equation of the line. in which case you'd get y+2 = (1/2)(x-3)

OpenStudy (anonymous):

ok thanks a bunchh! this stuff is really confusing me and i have a test tomm so i need to figure all this out

OpenStudy (anonymous):

not a problem at all! glad I could help!